# A Survey on Text-Driven 360-Degree Panorama Generation

### v2

@@ VI Challenges and Future Directions
Despite the impressive results achieved in text-driven 360-degree panorama generation, challenges remain in evaluation metrics, resolution, controllability, and model design. This section identifies these challenges and outlines potential directions for future research.
@@
Although OmniFID [51] extends FID to account for geometric fidelity in generated 360-degree panoramas, it primarily focuses on geometric aspects, leaving other crucial dimensions unaddressed. Text-to-image consistency, commonly evaluated by using the CLIP Score [52], is another critical metric for assessing text-driven 360-degree panorama generation methods. However, the CLIP Score, primarily designed for conventional 2D images, relies on an image encoder predominantly trained on standard perspective images. This encoder may not fully capture the inherent distortions and spatial relationships unique to 360-degree panoramas represented by equirectangular projections. Consequently, the CLIP Score may not accurately reflect the true semantic consistency in the context of 360-degree panorama generation. Future research could explore innovative evaluation metrics that effectively assess the semantic consistency between the input prompt and the generated 360-degree panoramic image.
@@
While Diffusion360 [58], which uses a super-resolution module, is among the few methods that can currently achieve a maximum resolution of 61443072 (6K), this remains inadequate for demanding applications like VR gaming and high-fidelity 3D scene reconstruction, which often necessitate resolutions of 8K or higher to capture intricate details of landscapes and architecture. Generating such high-resolution panoramas, however, introduces significant challenges related to computational latency and memory consumption. Addressing these limitations requires the development of more efficient model architectures and optimization techniques. Promising approaches include the use of window-based operations, model pruning, quantization, and advanced neural network designs tailored to resource-intensive tasks. Moreover, the availability of high-resolution, large-scale datasets will be critical for driving progress in this direction.
@@ VII Conclusion
This survey has provided a comprehensive overview of the rapidly evolving field of text-driven 360-degree panorama generation. We began by introducing primary representation methods of 360-degree panoramic images, along with widely used datasets, and key evaluation metrics in this domain. Subsequently, we presented an in-depth discussion of prevalent methods for text-driven 360-degree panorama generation and their applications in generating 360-degree 3D scene. Despite the significant progress achieved in this field, several challenges remain. To address these challenges, we have articulated promising directions for future research.

### v3

@@ VI Challenges and Future Directions
Despite the impressive results achieved in text-driven 360-degree panorama generation, challenges remain in evaluation metrics, resolution, controllability, model design, societal impact and industrial adoption. This section identifies these challenges and outlines potential directions for future research.
@@
As established in our analysis (see Sec. III-C), the development of metrics for global scene consistency remains a key challenge. This includes creating more robust, panorama-aware methods for evaluating both text-to-image semantic alignment and the overall structural plausibility of the generated 360-degree panoramic space. Future work could explore using advanced Vision-Language Models (VLMs) for question-based evaluations of complex spatial relationships. Another promising direction is the development of metrics that assess the implicit 3D geometry of the scene to detect logical inconsistencies in scale and layout that 2D metrics currently miss.
@@
While Diffusion360 [58], which uses a super-resolution module, is among the few methods that can currently achieve a maximum resolution of 61443072 (6K), this remains inadequate for demanding applications like VR gaming and high-fidelity 3D scene reconstruction, which often necessitate resolutions of 8K or higher to capture intricate details of landscapes and architecture. However, generating such high-resolution panoramas incurs both high memory consumption and long inference times, which severely limit practical deployment. Addressing these limitations will require the development of more efficient model architectures and optimization techniques. Promising approaches include the use of window-based operations, model pruning, quantization, knowledge distillation, and advanced neural network designs tailored to resource-intensive tasks. Moreover, the availability of high-resolution, large-scale datasets will be critical for driving progress in this direction.
@@
Ethical and Societal Considerations
The persuasive nature of 360-degree panoramas introduces risks of misuse, including fabricated environments, disinformation, and privacy violations. To mitigate these concerns, future research should pair technical advances with safeguards such as transparent data documentation, responsible licensing, and watermarking mechanisms. Proactive engagement with ethical guidelines and interdisciplinary oversight is essential to ensure that text-driven 360-degree panorama generation benefits society while limiting potential harm.
Industry Translation and Adoption
Text-driven 360-degree panorama generation is beginning to move from research to practice, supporting VR/AR content creation, virtual tourism, and game development. Major industry players, including Meta (Quest), Google (Street View), and Apple (Vision Pro), are actively developing related capabilities. Broader deployment is currently limited by requirements for stable 8K plus resolution, low latency on-device inference, and industry-standard benchmarks. Addressing these challenges through more efficient model design and closer collaboration between academia and industry will be critical for production-ready adoption.
@@ VII Conclusion
This survey has provided a comprehensive overview of the rapidly evolving field of text-driven 360-degree panoramic image generation. We began by introducing primary representation methods of 360-degree panoramic images, along with widely used datasets, and key evaluation metrics in this domain. Subsequently, we presented an in-depth discussion of prevalent methods for text-driven 360-degree panorama generation, and explored its two closely related directions: text-driven 360-degree 3D scene generation and text-driven 360-degree panoramic video synthesis. Despite the significant progress achieved in this field, several challenges remain. To address these challenges, we have articulated promising directions for future research.